OpenStudy (anonymous):

Does this have any symmetry? (2x^2 -5x + 2)/(x^2 -4) It asked if it had even, odd, or no symmetry. I said no.

OpenStudy (amistre64):

replace x with -x f(-x) = f(x) ... its even f(-x) = -f(x) its odd anything else is neither

OpenStudy (amistre64):

even exponents give back positive x's odd exponents give back negative exponents 2x^2 +5x +2 ------------ not= f(x); not= -f(x) ... i say neither x^2 -4

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=%282x^2+-5x+%2B+2%29%2F%28x^2+-4%29 the graph confirms it; it aint mirrored over the y or the origin ...